OpenStudy (anonymous):

Series and sequences question Write the first five terms of the sequence: 1. a1=3, a2=-2 and an+2=(an+1)/an for n greater than or equal to 1. ....meh.

OpenStudy (anonymous):

You just substitute in the values: a3 = a(n+2) = a(1+2) = (a1+1)/a(1) = a2/a1 = -2/3. Now you have a3. Calculate a4. Then keep going.

OpenStudy (anonymous):

You probably want to double-check these, it's getting late.....\[a_1=3,~ a_2=-2 \implies a_3=-2/3\implies a_4=1/3\implies a_5=-1/2\]

OpenStudy (anonymous):

Thanks guys! But alas, I see how a3 is found. a4 and a5 allude me.

OpenStudy (anonymous):

\[a_4=\frac{a_3}{a_2}\]

OpenStudy (anonymous):

\[a_5=\frac{a_4}{a_3}\]

OpenStudy (anonymous):

Got it now?

OpenStudy (anonymous):

:D YEP!

OpenStudy (anonymous):

Gotta love it when that happens....

OpenStudy (anonymous):

I love those smacks in the face! HAHA i worked 9 hours today and it is late :P